A) A 1.00 g gaseous sample of hydrocarbon occupies a volume of 385 mL at 330 K and 1.00 atm. Find the molar mass of the compound.

B) Calculate the number of atoms of He(g) that are needed to exert a pressure of 4.0 X 10–2 kPa in a 15 mL container at a temperature of –23°C.


1
Expert's answer
2021-01-25T04:21:07-0500

A) A 1.00 g gaseous sample of hydrocarbon occupies a volume of 385 mL at 330 K and 1.00 atm. The molar mass of the compound is found using gas law:

p*V=m/M*RT

M = m*RT/(p*V) = 1 g * 8.31 J/(K*mole)*330 K / (385*10-6 m3 * 101325 Pa) = 70.30 g/mole


B) The number (N) of atoms of He(g) that are needed to exert a pressure of 4.0 X 10–2 kPa in a 15 mL container at a temperature of –23°C is found using gas law:

p*V = N*k*T

N = p*V/(k*T) = 40 Pa * 15*10-6 m3 / (1.38*10-23 J/K * 350 K) = 1.27*1017